A ball at the end of a string is revolved at a uniform rate in a vertical circle of radius 72.0 cm. If its speed is 4.00m/s and its mass is 0.30 kg calculate the tension in the string when the ball is at the top and bottom. Include equations and solutions
1
Expert's answer
2020-05-03T17:31:33-0400

Top:


"T_t=mg+m\frac{v^2}{r}\\=(0.3)(9.8)+0.3\frac{4^2}{0.72}=9.6\ N"

Bottom:


"T_b=-mg+m\frac{v^2}{r}\\=-(0.3)(9.8)+0.3\frac{4^2}{0.72}=3.7\ N"